OpenStudy (anonymous):

can any one help me check my answers for a couple of problems. medal and fan rewarded!!

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6. y = one thirdx + 14 thirds y = 3x + 10 y = -3x - 2 y = -one thirdx + 10 thirds

OpenStudy (anonymous):

for this one i picked y = -3x - 2

OpenStudy (johnweldon1993):

That would be correct!

OpenStudy (anonymous):

can you help me on more??

OpenStudy (johnweldon1993):

Sure, go ahead and post em

OpenStudy (anonymous):

ok ill tag you in another one but its about 8 questions are you ok with that?